OpenStudy (anonymous):

If rate constant for a 1st order reaction is k, then T1/4 can be written as: 0.29/k. I know, that k=0.693/T1/2, where k=> rate const. and T1/2=> half life. T1/4 is twice half-life. How do you get the given answer?i.e 0.29/k.

OpenStudy (anonymous):

T (1/4) = [ln (1/(1-0.25))] /k = ln (4/3) x 1/k = 0.285/k

OpenStudy (anonymous):

How can i use T1/2 = 0.693/k, without using logs.

OpenStudy (anonymous):

uve got to use logs even if u break it up like [ln 4 - ln 3]/k = 2(ln2)/k -(ln3)/k the first term u cn use t1/2 but uve gotta use logs for the second term